### v2

# A weak regularity lemma for polynomials

Abstract
A regularity lemma for polynomials provides a decomposition in terms of a bounded number of approximately independent polynomials. Such regularity lemmas play an important role in numerous results, yet suffer from the familiar shortcoming of having tower-type bounds or worse. In this paper we design a new, weaker regularity lemma with strong bounds. The new regularity lemma in particular provides means to quantitatively study the curves contained in the image of a polynomial map, which is beyond the reach of standard methods.
Applications include strong bounds for a problem of Karam on generalized rank, as well as a new method to obtain upper bounds for fan-in parameters in arithmetic circuits. For example, we show that if the image of a polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ of degree $d$ does not contain a line, then $\mathbf{P}$ can be computed by a depth-$4$ arithmetic formula with bottom fan-in at most $d/2$ and top fan-in at most $(2m)^{C(d)}$ (with $C(d)=2^{(1+o(1))d}$). One implication of our work is a certain “barrier” to arithmetic circuit lower bounds, in terms of the smallest degree of a polynomial curve contained in the image of the given polynomial map.
1 Introduction
Regularity lemmas for polynomials have served as a key component in a variety of important results, as in the proof of the the finite-field Gowers Inverse conjecture in additive combinatorics and higher-order Fourier analysis (e.g. [GT09, TZ12]), Stillman’s conjecture in commutative algebra [AH20], and the Reed-Muller weight distribution problem in coding theory [BL15, KLP12], to name but a few. They serve as a key component in higher-order arithmetic regularity lemmas (e.g., [GW11, Gree06, GT09, HL11, KL08]). [^1] Roughly speaking, arithmetic regularity lemmas approximate general functions in terms of low-degree polynomials that are approximately independent. (This approximate independence is indispensable; to quote Green [Gree06]: second-order arithmetic regularity lemmas “reduce the study of general functions… to the study of projections… onto quadratic factors. This, however, is of little consequence unless we can study those supposedly simple objects.”) As is usually the case with regularity lemmas, they unfortunately only have weak bounds, and therefore so does any result that relies on them. In this paper we introduce a weaker regularity lemma for polynomials, with far better bounds, that is nevertheless powerful enough for some applications. For example, we obtain strong bounds for a result of Karam [Kar23] on the relation between the generalized rank of a polynomial and its range. Our results also yield a method to prove upper bounds for arithmetic circuits.
Roughly speaking, the usual regularity lemmas for polynomials (see, e.g., Proposition 8.1 in [ESS19], Lemma 2.4 in [GT09], or Lemma 3.11 in [Gree06]) can be described as results about polynomial decompositions. Namely, they express any $n$-variate polynomial $P\colon\mathbb{F}^{n}\to\mathbb{F}$—and more generally, any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$—as in terms of a bounded number (independent of $n$) of polynomials $\mathbf{X}=(X_{1},...,X_{k})$, that is, $P=F(\mathbf{X})$, with the property that the polynomials in $\mathbf{X}$ “behave” approximately like $k$ (jointly independent) variables. Although the precise definition of “behave” can differ depending on the context, the intuition is as follows: for many purposes, any polynomial can be analyzed as though it had only $k$ variables. The number $k$ of parts $X_{1},\ldots,X_{k}$ in the decomposition depends only on the degree $d$ and the number $m$ of the given polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$, and as mentioned above, this dependency—befitting a regularity lemma—can be quite astronomical, often having worse than a tower-type growth (see, e.g., the Remark following Lemma 2.4 in [GT09]). [^2] Interestingly, recent regularization results of Lampert and Ziegler [LZ24, Lam23] give strong bounds; however, unlike regularity lemmas, these do not provide a decomposition of $\mathbf{P}$ as a polynomial in $X_{1},\ldots,X_{k}$, but rather place $\mathbf{P}$ in the ideal generated by them. The main reason for these weak bounds is that in order to guarantee approximate independence, the polynomial parts $X_{1},\ldots,X_{k}$, as well as all their linear combinations, are required to have rank (see Section 2) that is large as a function of $k$, their number. Thus, the larger the number of parts, the more difficult it is to satisfy the rank requirement, which translates to an extremely large number of steps in the process to construct them. [^3] See, e.g., Proposition 8.1 in [ESS19] for a clean example of this phenomenon, where the only thing that guarantees that the process at all terminates is a small decrease in the lexicographic order of the tuple of degrees of $\mathbf{X}$ in each step.
In the setting of graphs, Frieze and Kannan [FK99] famously obtained a weak regularity lemma that has both strong bounds and multiple interesting applications. For comparison, the usual graph regularity lemma of Szemerédi decomposes any graph—by partitioning its vertices—into a bounded number of parts (independent of the number of vertices) with the guarantee that almost all graph parts are “approximately” random. The weak graph regularity lemma decomposes any graph in the same sense, but comes with only one, global guarantee. Our weak regularity lemma for polynomials is partly inspired by this approach.
1.1 Weak regularity for polynomials
The basic idea of our weak regularity lemma is to relax the requirement on the $k$ polynomials $X_{1},\ldots,X_{k}$, and only require that at least one of them, of maximal degree, behaves like a free variable with respect to all the others. That is, conditioned on $X_{2},\ldots,X_{k}$ taking any given value $y\in\mathbb{F}^{k-1}$, we require that $X_{1}$ takes any value in $\mathbb{F}$ with approximately the same frequency. This imitates the case of a genuinely free variable; or more concretely, the case where $X_{1}=x_{1}$, and $X_{i}=X_{i}(x_{2},\ldots,x_{n})$ for $i>1$, where $x_{1},\ldots,x_{n}$ are variables.
Intuitively, the benefit of having a weak regular decomposition for a polynomial map $\mathbf{P}$ is that it reveals some information about the univariate polynomials “induced” by $\mathbf{P}$. In particular, it enables us to argue about the image of $\mathbf{P}$ (for which standard rank-based tools are unsuitable, as they typically give information only about equidistribution); indeed, a corollary of the weak regularity lemma will be that any line that $\mathbf{X}$ intersects somewhat often must in fact be fully contained in the image of $\mathbf{X}$. This translates to a polynomial curve (univariate polynomial map) contained in the image of $\mathbf{P}$. See below for more about applications.
We note that throughout the paper we work with finite fields $\mathbb{F}$, for which the notion of weak regularity can be most elementarily defined. Moreover, rank-based methods, which we rely on extensively in the proof (as do proofs of other regularity lemmas for polynomials), tend to be most challenging in the finite-field setting. [^4] To get an analogue of our weak regularity lemma for infinite fields would require replacing the probabilistic notion of approximate independence in the definition of weak regularity below, for example by a geometric or by a commutative algebra analogue (see, e.g., [LZ24, ESS19]), and adapting the proof accordingly. Technically, this would require replacing the notion of bias of a polynomial (see Section 2.2) by standard analogues for infinite fields. For the structure-vs-randomness theorem we rely on (Theorem 2.10), a geometric analogue over infinite fields has long been known [KLP24, Sch85] with, in fact, a slightly better, linear bound.
Formally, we define weak regularity as follows. We refer the reader to Section 2 for a more formal discussion, and the proofs. For tuples of polynomials $\mathbf{P}=(P_{1},\ldots,P_{m})$ and $\mathbf{X}=(X_{1},\ldots,X_{k})$ over $\mathbb{F}$, we write $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ if every component $P_{i}$ of $\mathbf{P}$ can be written as $F_{i}(X_{1},\ldots,X_{k})$ for some polynomial $F_{i}$. Henceforth, for any $k$-tuple $v=(v_{1},\ldots,v_{k})$, we use the notation $v^{\prime}$ for the (k1)(k-1)-tuple $v^{\prime}=(v_{2},\ldots,v_{k})$.
Definition (weak regularity).
A tuple of polynomials $\mathbf{P}$ over $\mathbb{F}=\mathbb{F}_{q}$ has a weak $\epsilon$-regular decomposition of size $k$ if $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$, for some $k$ homogeneous polynomials $\mathbf{X}=(X_{1},\ldots,X_{k})$, such that:
  • for every $y\in\mathbb{F}^{k}$, $|\Pr[\mathbf{X}=y]-q^{-1}\Pr[\mathbf{X}^{\prime}=y^{\prime}]|\leq\epsilon q^{-k}$,
  • $\mathbf{P}\nsubseteq\mathbb{F}[\mathbf{X}^{\prime}]$, and
  • $\deg(X_{1})=\max_{i}\deg(X_{i})$.
The probabilities here are with respect to a uniformly random input; that is, $\Pr[\mathbf{X}=y]$ is a shorthand for $\Pr_{x\in\mathbb{F}^{n}}[\mathbf{X}(x)=y]$, with $x$ chosen uniformly at random from $\mathbb{F}^{n}$. Intuitively, the probabilistic condition can be motivated as follows. If $X_{1}$ were truly an independent variable $x_{1}$, with $X_{i}=X_{i}(x_{2},\ldots,x_{n})$ for $i>1$, we would certainly have both $\Pr[X_{1}=y_{1}]=\frac{1}{q}$ and $\Pr[\mathbf{X}=y]=\Pr[X_{1}=y_{1}]\Pr[\mathbf{X}^{\prime}=y^{\prime}]$. Hence, the probabilistic condition mimics the behavior we would expect for an independent variable $x_{1}$.
Note that the two additional conditions at the end of the definition are necessary to avoid trivialities. This is clearly the case for the first, which guarantees that the decomposition of $\mathbf{P}$ depends on $X_{1}$. As for the other additional condition, that $X_{1}$ is of maximal degree—note that we can decompose any polynomial $P=P(x_{1},\ldots,x_{n})$, of degree $d$ in $x_{1}$ say, as $\sum_{i=0}^{d}x_{1}^{i}C_{i}(x_{2},\ldots,x_{n})$; that is, we always have a small decomposition $P=F(x_{1},C_{0},\ldots,C_{d})$ (where $F(z,y_{0},y_{1},\ldots,y_{d})=\sum_{i=0}^{d}z^{i}y_{i}$). Although $x_{1}$ is genuinely independent of the other parts of the decomposition $C_{0},\ldots,C_{d}$, its degree $\deg(x_{1})=1$ is generally far from the largest degree among them.
Let us also mention two extreme cases in the definition of weak regularity. First, if the polynomial tuple $\mathbf{P}$, viewed as a map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$, is itself equidistributed, that is, $|\Pr[\mathbf{P}=y]-q^{-1}|\leq\epsilon q^{-m}$ for every $y\in\mathbb{F}^{m}$, then $\mathbf{P}$ trivially has a small weak $\epsilon$-regular decomposition (with $\mathbf{X}=\mathbf{P}$). Second, $\mathbf{P}$ always has the trivial decomposition $\mathbf{P}=\mathbf{P}(\mathbf{X})$ with $\mathbf{X}=(x_{1},\ldots,x_{n})$ its variables, which is a weak 0-regular (i.e., a perfectly weak regular decomposition; assuming without loss of generality that $\mathbf{P}$ depends on $x_{1}$), though of a trivially large size $n$.
Our main result is that a rather small weak regular decomposition always exists (see Theorem 2.2).
Theorem (weak regularity lemma).
Let $\epsilon=q^{-r}$ with $r>1$. Every polynomial $m$-tuple $\mathbf{P}$ over $\mathbb{F}=\mathbb{F}_{q}$, of degrees at most $d$ with $d<\operatorname{char}(\mathbb{F})$, has a weak $\epsilon$-regular decomposition of size at most $(mr)^{2^{d(1+o(1))}}$.
For comparison, the usual regularity lemmas for polynomials only give tower-type bounds or worse; namely, the bounds are at the very least a tower of exponentials, of height depending on $d$, with $m$ at the top. Our weak regularity lemma, in particular, is polynomial in $m$ (which is especially relevant in the important case of low-degree polynomials $d=O(1)$).
1.2 Applications
Is there an algebraic explanation for non-surjectivity of (multivariate) polynomials? An important class of non-surjective polynomials are those of the form $P=F(Q)$ with $1<\deg(Q)<\deg(P)$—sometimes called uni-multivariate decomposable [GRS02]—for some non-surjective univariate $F$ (for example, $P=(Q(x_{1},...,x_{n}))^{t}$ with $\gcd(t,|\mathbb{F}|-1)>1$). For such a polynomial of degree $d$, it must be that $\deg(Q)\leq d/2$ (as $\deg(F)\geq 2$). We show that this is always the case, in the sense that every non-surjective polynomial can be written in terms of a rather small number of polynomials all of degree at most $d/2$ (despite most polynomials not being uni-multivariate decomposable [VZ15]). In fact, we show a significant generalization of this to polynomial maps. Note that the image of a polynomial map holds information about the dependencies between its various polynomials.
Following Green and Tao ([GT09], Definition 1.6), let $\operatorname{rank}_{t}(P)$, for a polynomial $P$ and $t\geq 0$, denote the least number of polynomials of degree at most $t$ such that $P$ can be expressed as a function of only them. Note that this extends naturally to polynomial maps $\mathbf{P}$. One easy-to-state application of the weak regularity lemma is that any polynomial that is not surjective, and more generally any polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ whose image $\operatorname{Im}\mathbf{P}$ ($\subseteq\mathbb{F}^{m}$) does not contain a line, must have low generalized rank $\operatorname{rank}_{t}(\mathbf{P})$, for some $t$ bounded away from $\deg(\mathbf{P}):=\max_{i}\deg(P_{i})$.
Theorem (see Theorem 3.3).
Let $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ be a polynomial map with $\deg(\mathbf{P})\leq d$, over any finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. If $\operatorname{Im}\mathbf{P}$ does not contain a line then
$\operatorname{rank}_{d/2}(\mathbf{P})\leq(2m)^{2^{d(1+o(1))}}.$
One can significantly extend this result, by considering the largest degree of a curve contained in the image of $\mathbf{P}$. Formally, we define the univariate degree of a polynomial map $\mathbf{P}$ as
$\operatorname{udeg}(\mathbf{P})=\min\{\deg(\mathbf{U})\,|\,\mathbf{U}\colon\mathbb{F}\to\mathbb{F}^{m}\text{ non-constant}\colon\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}\}.$
We note that $\operatorname{udeg}(P)\leq\deg(P)$, and that $\operatorname{udeg}(\mathbf{P})=1$ if and only if $\operatorname{Im}\mathbf{P}$ contains a line; see Section 3.1 for more properties of univariate degree. As an example, we have $\operatorname{udeg}(Q^{t})\leq t$, for any polynomial $Q$, with equality for any surjective $Q$ (in which case $\operatorname{Im}(Q^{t})=\{x^{t}|x\in\mathbb{F}\}$) and some choices of $t$ that depends on $|\mathbb{F}|$. Note that in this last example $P=Q^{t}$, we have $\deg(P)/\operatorname{udeg}(P)=\deg(Q)$, and if we take $Q$ to be a random polynomial $Q\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\operatorname{rk}_{\deg(Q)-1}(P)$ is unbounded (as $n\to\infty$; see Example 1.4 in [Kar23]). This shows that we cannot hope to bound $\operatorname{rk}_{t}(P)$ for $t<\deg(P)/\operatorname{udeg}(P)$.
Karam [Kar23] proved that $\operatorname{rank}_{t}(P)$, for $t=\deg(P)/\operatorname{udeg}(P)$, is, in fact, bounded (independently of nn). [^5] Karam’s result actually bounds the closely related notion of $\operatorname{rk}_{t}$; see Section 3 below.
Theorem 1.1 (Karam, Theorem 1.3 in [Kar23]).
Let $P\colon\mathbb{F}^{n}\to\mathbb{F}$ be a polynomial with $\deg(P)\leq d$, over any prime finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. For $t=\deg(P)/\operatorname{udeg}(P)$, we have $\operatorname{rank}_{t}(P)\leq\gamma(p,d)$, for some function $\gamma$.
The proof in [Kar23] uses a variant of the regularity lemma for polynomials, and thus yields an extremely weak bound $\gamma(p,d)$ on $\operatorname{rk}_{t}(P)$.
We show that the weak regularity lemma is sufficiently powerful to bound $\operatorname{rank}_{t}$ for any polynomial, and in fact also for any polynomial map, yielding a correspondingly strong bound.
Theorem (see Theorem 3.7).
Let $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ be a polynomial map with $\deg(\mathbf{P})\leq d$, over any finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. For $t=\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})$,
$\operatorname{rank}_{t}(\mathbf{P})\leq(2m)^{2^{d(1+o(1))}}.$
Roughly speaking, the proof follows by showing that for any weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, for $\epsilon$ bounded away from $1$, we have $\operatorname{udeg}(\mathbf{P})\leq\deg_{1}(\mathbf{F})$, the degree of $\mathbf{F}$ in its first variable (so $\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}$ for some non-constant curve $\mathbf{U}$ with $\deg(\mathbf{U})=\deg_{1}(\mathbf{F})$). Using the guarantees of the homogeneity of $\mathbf{X}$ and of the maximality of $\deg(X_{1})$ ($\deg(\mathbf{X})$), we will deduce $\deg_{1}(\mathbf{F})\leq\deg(\mathbf{P})/\deg(\mathbf{X})$, that is, $\deg(\mathbf{X})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})=t$, which would allow us to bound $\operatorname{rk}_{t}$.
Finally, this machinery also has a direct implication to the study of arithmetic formulas in computational complexity theory, namely, depth-$4$ arithmetic formulas. A $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula (e.g., [KSS14, KS14]) for a polynomial $P$ is a decomposition of the form $P=\sum_{i=1}^{r}\prod_{j=1}^{a_{i}}Q_{i,j}$ with $a_{i}\leq a$ and $\deg(Q_{i,j})\leq b$. It is an important challenge to understand the interplay between any of the different “fan-in” parameters $(r,a,b)$. Trivially, if the bottom fan-in $b$ is allowed to be as large as $\deg(P)$, then $P$ has a $\sum^{[r]}\prod^{[a]}\sum\prod^{[b]}$-formula with $r=a=1$. The question becomes interesting when the bottom fan-in $b$ is bounded away from $\deg(P)$. Using the weak regularity lemma, we deduce an upper bound on the top fan-in $r$ from a lower bound on the univariate degree. [^6] It is worth noting that the model of depth-$4$ homogeneous arithmetic formulas with top fan-in $r$ is already powerful for $r=\Theta(\log n)$; see [KS14].
Theorem 1.2.
Let $P\colon\mathbb{F}^{n}\to\mathbb{F}$ with $\deg(P)\leq d$, over any finite field $\mathbb{F}$ with $d<\operatorname{char}(\mathbb{F})$. Suppose $P\in\mathbb{F}[\mathbf{P}]$ with $\mathbf{P}=(P_{1},\ldots,P_{m})$, $\deg(\mathbf{P})\leq d$, such that $\operatorname{udeg}(\mathbf{P})\geq u$. Then there is a $\sum^{[r]}\prod^{[2u]}\sum\prod^{[d/u]}$-formula for $P$ with top fan-in $r\leq(2m)^{2^{d(1+o(1))}}$. Moreover, the formula is homogeneous [^7] That is, each $Q_{i,j}$ is homogeneous, and all products $\prod_{j}Q_{i,j}$ are of the same degree. if $P$ is.
Contrapositively, this means that any strong enough lower bound method for the top fan-in of depth-$4$ arithmetic formulas (of a given bottom fan-in, size, etc.) must somehow avoid applying to polynomial maps of high univariate degree, and to the polynomials they generate. It therefore seems that the notion of univariate degree could be an important parameter to consider when studying questions about arithmetic circuits and their top fan-in, and for polynomial identity testing of depth-$4$ formulas with bounded fan-in (a topic of considerable interest, see e.g., [DDS21]).
Remark 1.3.
Univariate degree should not be confused with the concept of an elusive function of Raz [Raz10], which is a polynomial map $\mathbf{P}$ satisfying $\operatorname{Im}\mathbf{P}\nsubseteq\operatorname{Im}\mathbf{\Gamma}$, for every polynomial map $\mathbf{\Gamma}$ with given parameters. Indeed, univariate degree is about containment in the other direction: a lower bound on $\operatorname{udeg}(\mathbf{P})$ means that $\operatorname{Im}\mathbf{U}\nsubseteq\operatorname{Im}\mathbf{P}$, for every polynomial curve $\mathbf{U}$ of a given degree.
1.3 The proof
Towards the proof of the weak regularity lemma, we obtain results on linear spaces of polynomials and forms (homogeneous polynomials) that may be of independent interest. In particular, we analyze subspaces of high rank polynomials (where the rank is allowed to grow with the number of spanning polynomials, i.e., the dimension of the space). More particularly, we analyze high-rank polynomial “pencils”. By a polynomial pencil we mean a set of the form $V\setminus U$ where $V$ is a linear space of polynomials and $U$ is a strict subspace. For comparison, the classical regularity lemmas for polynomials work only with the “trivial” pencil $V\setminus\{0\}$; or more concretely, all nonzero linear combinations of the parts $\mathbf{X}$ are required to have high rank. Obtaining a high-rank pencil, rather than a high-rank trivial pencil, turns out to be achievable with far better bounds. We call this result a “rank-regularity lemma”, and note that its strong bounds eventually translate to strong bounds for our weak regularity lemma. Using it, the rest of our proof has, very roughly, the following logical structure (see Section 2 for all definitions):
$\mathbf{X}\text{ $t$-rank-regular}$ $\Leftrightarrow\operatorname{rk}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})>t|\mathbf{X}|$
$\Rightarrow|\operatorname{bias}(X_{1}+\operatorname{Sp}\mathbf{X}^{\prime})|<\epsilon q^{-|\mathbf{X}|}$
$\Rightarrow\mathbf{X}\text{ weak $\epsilon$-regular.}$
The proof moreover guarantees the additional required conditions, e.g., that $X_{1}$ has maximal degree. To translate from high rank to low bias as above, our proof relies on recent results on the relationship between structure and randomness for polynomials (see e.g. [Jan20, Mil19, CM21, MZ24]). Incidentally, these results are proved by going through the corresponding multilinear forms—which requires the characteristic of the field to be larger than the degree of the parts—by relating analytic rank and partition rank (defined by Gowers and Wolf [GW11] and by Naslund [Nas20], respectively).
1.4 Organization
In Section 2 we state and prove our weak regularity lemma. The proof has several components. In Section 2.1 we prove a rank-regularity lemma, which finds a small decompositions where every linear combination outside of a strict subspace has high rank. In Section 2.2 we show that a similar condition in terms of bias (rather than rank) suffices to get weak regularity, using Fourier-analytic techniques and known structure-vs-randomness results. Finally, in Section 2.3, we combine the aforementioned pieces to prove the weak regularity lemma. In Section 3 we obtain the applications of the weak regularity lemma outlined above.
1.5 Preliminaries
An important aspect of regularity lemmas for polynomials is that they are independent of the number of variables $n$ (similarly to how, for example, the graph regularity lemma is independent of the number of vertices of the graph). As such, it will be convenient to use the notation $\operatorname{Poly}(\mathbb{F})$ for the set of all polynomial functions $P\colon\mathbb{F}^{n}\to\mathbb{F}$ (or polynomials for short) over $\mathbb{F}$ in any (finite) number $n$ of variables. Note that any $P\in\operatorname{Poly}(\mathbb{F})$ has a unique representation as (and so can be identified with) an $\mathbb{F}$-linear combination of monomials with each variable of degree at most $|\mathbb{F}|-1$; in particular, the (total) degree $\deg(P)$ is well defined, as well as the support $\operatorname{Supp}(P)$. We write $\operatorname{Poly}_{d}(\mathbb{F})$ for the set of polynomials of degree at most $d$. We henceforth omit $\mathbb{F}$ from our notation when it is clear from context. Naturally, $\operatorname{Poly}^{m}=\{(P_{1}\ldots,P_{m})|\forall i\colon P_{i}\in\operatorname{Poly}\}$ stands for the set of $m$-tuples of polynomials, or equivalently, polynomial maps into $\mathbb{F}^{m}$; and analogously for $\operatorname{Poly}_{d}^{m}$. Similarly, we denote by $\operatorname{Form}$ the set of all homogeneous polynomials (or forms for short) of positive degree, in any (finite) number of variables; we denote by $\operatorname{Form}_{d}$ the set of forms of degree $d$ ($\geq 1$); and by $\operatorname{Form}_{d}^{m}$ the set of $m$-tuples of forms of degree $d$. For polynomials $X_{1},\ldots,X_{r}\in\operatorname{Poly}(\mathbb{F})$, we denote by $\mathbb{F}[X_{1},\ldots,X_{r}]$ the algebra they generate, that is, $\{F(X_{1},\ldots,X_{r})|r\text{-variate }F\in\operatorname{Poly}\}$. Of course, if $x_{1},\ldots,x_{n}$ are independent variables then $\mathbb{F}[x_{1},\ldots,x_{n}]$ is the usual ring of $n$-variate polynomials in $x_{1},\ldots,x_{n}$.
We will frequently use boldface letters to emphasize that an object is a tuple; so for example, we will use $\mathbf{P}$, instead of $P$, when dealing with a tuple of polynomials (or a polynomial map). We will use the convention that for a tuple $\mathbf{X}=(X_{1},\ldots,X_{k})$ and a set $A$, the notation $\mathbf{X}\subseteq A$ is a shorthand for $X_{i}\in A$ for all $i$. For $\mathbf{P}\subseteq\operatorname{Poly}$ we denote $\deg(\mathbf{P})=\max_{i}P_{i}$. More generally, for a set $V\subseteq\operatorname{Poly}$ we denote $\deg(V)=\max_{P\in V}\deg(P)$.
Probabilities, unless otherwise specified, are to be understood with respect to a uniformly random input; that is, for $\mathbf{X}\colon\mathbb{F}^{n}\to\mathbb{F}^{k}$, $\Pr[\mathbf{X}=y]$ is a shorthand for $\Pr_{x\in\mathbb{F}^{n}}[\mathbf{X}(x)=y]$ with $x$ chosen uniformly at random from $\mathbb{F}^{n}$. The image (or range) of a polynomial map $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ is denoted $\operatorname{Im}\mathbf{P}=\{\mathbf{P}(x)|x\in\mathbb{F}^{n}\}$ ($\subseteq\mathbb{F}^{m}$). For a subset $S\subseteq\mathbb{F}^{n}$, we write $\mathbf{P}(S)=\{\mathbf{P}(x)|x\in S\}$.
Note that $\operatorname{Poly}_{d}(\mathbb{F})$ is a linear space over $\mathbb{F}$. We write $U\leq V$ when $U$ is a subspace of $V$. We denote the linear span of $\mathbf{X}\in\operatorname{Poly}_{d}^{k}(\mathbb{F})$ by $\operatorname{Sp}\mathbf{X}=\operatorname{Sp}\{X_{1},\ldots,X_{k}\}=\{a\cdot\mathbf{X}|a\in\mathbb{F}^{k}\}$, which is of course a subspace of $\operatorname{Poly}_{d}(\mathbb{F})$.
For $\mathbf{P}\in\operatorname{Poly}^{m}(\mathbb{F})$, we refer to an expression of the form $\mathbf{P}=\mathbf{F}(\mathbf{X})$ (or $\mathbf{P}=\mathbf{F}\circ\mathbf{X}$), with $\mathbf{F}\in\operatorname{Poly}^{m}(\mathbb{F})$ and $\mathbf{X}\in\operatorname{Poly}^{k}(\mathbb{F})$, as a decomposition of $\mathbf{P}$ (so that for every component $P_{i}$ of $\mathbf{P}$, we have the decomposition $P_{i}=F_{i}(\mathbf{X})$). When the map $\mathbf{F}$ is immaterial, we denote the decomposition by $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$.
As mentioned above, all fields considered in this paper are finite; we denote by $\mathbb{F}_{q}$ the finite field of size $q$ (so $q$ is a power of the characteristic $\operatorname{char}(\mathbb{F})$).
Acknowledgments:
GM was supported by NSF Award DMS-2302988 and a PSC-CUNY award.
2 Weak Regularity
For convenience, we repeat here the definitions of weak regularity and the weak regularity lemma. Recall that for a $k$-tuple $\mathbf{y}=(y_{1},\ldots,y_{k})$, we denote by $\mathbf{y}^{\prime}$ the (k1)(k-1)-tuple $\mathbf{y}^{\prime}=(y_{2},\ldots,y_{k})$.
Definition 2.1 (weak $\epsilon$-regularity).
Let $\mathbb{F}=\mathbb{F}_{q}$. We say that $\mathbf{X}\in\operatorname{Form}^{k}(\mathbb{F})$ is weak $\epsilon$-regular if $\deg(X_{1})=\deg(\mathbf{X})$, and for every $y\in\mathbb{F}^{k}$,
$\big|\Pr[\mathbf{X}=\mathbf{y}]-q^{-1}\Pr[\mathbf{X}^{\prime}=\mathbf{y}^{\prime}]\big|\leq\epsilon q^{-k}.$
For $\mathbf{P}\subseteq\operatorname{Poly}(\mathbb{F})$, we say that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ is a weak $\epsilon$-regular decomposition if $\mathbf{X}\subseteq\operatorname{Form}$ is weak $\epsilon$-regular and $\mathbf{P}\nsubseteq\mathbb{F}[\mathbf{X}^{\prime}]$.
Roughly speaking, a set of forms $\mathbf{X}\subseteq\operatorname{Form}$ is weak $\epsilon$-regular if at least one of its members, of largest degree, is approximately “unbiased” and “independent” of the rest; moreover, a polynomial map $\mathbf{P}$ has a weak $\epsilon$-regular decomposition if it can be expressed in terms of a weak $\epsilon$-regular $\mathbf{X}$, and depends on $X_{1}$. Intuitively, $P$ can be thought of as a polynomial in only $k$ variables, and although these may not be independent, free variables, at least one of them approximately is.
Let us also note that the probabilistic condition in Definition 2.1 means that $\Pr[\mathbf{X}=\mathbf{y}]/q^{-k}=\alpha\pm\epsilon$ for every $\mathbf{y}\in\mathbb{F}^{k}$ with $\Pr[\mathbf{X}^{\prime}=\mathbf{y}^{\prime}]/q^{-(k-1)}=\alpha$ (recall $\mathbf{X}^{\prime}$ is (k1)(k-1)-tuple). Intuitively, this means $\mathbf{X}$ is distributed like $\mathbf{X}^{\prime}$ up to an additive error of $\epsilon$.
Theorem 2.2 (weak regularity lemma).
Let $\epsilon>0$. Every $\mathbf{P}\in\operatorname{Poly}_{d}^{m}(\mathbb{F})$, with $d<\operatorname{char}(\mathbb{F})$, has a weak $\epsilon$-regular decomposition, for $\epsilon=q^{-r}$, of size at most
$(2m(r+1))^{2^{d(1+o(1))}}.$
Note that the above bound is polynomial in $m$. In particular, if $\epsilon=\Omega(1)$ then the bound in Theorem 2.2 simplifies to $(2m)^{C}$, with $C\leq 2^{d(1+o(1))}$. On the other hand, if $\epsilon=q^{-r}$ with $r$ bounded away from 0, then the bound simplifies to $(2mr)^{C}$ (and to $(mr)^{C}$ if $r$ is bounded away from $1$). In fact, a slightly more careful analysis (see the proof in Section 2.3) gives this bound even when one replaces the $\epsilon q^{-k}$ error term in Definition 2.1 by $q^{-rk}$.
2.1 Rank-regularity
The proof of the weak regularity has several components. At the core of the proof is what we call a rank-regularity lemma [^8] Or “rankularity” lemma for short. (see Definition 2.4). Roughly, the requirement for the parts $\mathbf{X}$ in a rank-regular decomposition is that they are homogeneous (not necessarily all of the same degree), and that every linear combination of $\mathbf{X}$ outside of a strict subspace has high rank (see Definition 2.3). We will later convert a rank-regular decomposition into a weak regular decomposition. We begin by setting up the basic definitions we need, and state the results that form the steps of our overall strategy.
An important definition in the study of polynomials is that of rank, sometimes also called strength [AH20] or Schmidt rank [Sch85]. Recall that a polynomial is reducible if it can be factorized as the product of two non-constant polynomials (all polynomials over the same field).
Definition 2.3 (rank).
The rank of a form $P\in\operatorname{Form}_{d}(\mathbb{F})$ is:
  • For $d\geq 2$: the least number of reducible forms that sum up to $P$:
    $\operatorname{rk}(P)=\min\big\{r\,\big|\exists R_{1},\ldots,R_{r}\in\operatorname{Form}_{d}(\mathbb{F})\text{ reducible}\colon P=R_{1}+\cdots+R_{r}\big\}.$
  • For $d=1$, $\operatorname{rk}(P)=\infty$.
More generally, the rank of a polynomial $P\in\operatorname{Poly}(\mathbb{F})$ is the rank of its highest-degree portion, [^9] That is, if $P$ is of degree $d$ and we write $P=\sum_{i=0}^{d}P_{i}$ with $P_{i}$ a form of degree $i$, then $\operatorname{rk}(P)=\operatorname{rk}(P_{d})$. or 0 if $P$ is a constant.
Note that for a homogeneous $P\in\operatorname{Form}_{d}$ with $\operatorname{rk}(P)\leq r$ ($\infty$), we can write $P=Q(R_{1},\ldots,R_{k})$ with $k\leq 2r$, for some homogeneous $Q\in\operatorname{Form}_{2}$ and $R_{i}\in\operatorname{Form}$ with $\deg(R_{i})<d$.
The core of the proof of our weak regularity lemma is a weak regularity lemma for $\operatorname{rk}$. For $\mathbf{P}\subseteq\operatorname{Poly}$, call a decomposition $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ minimal if $\mathbf{P}\nsubseteq\mathbb{F}[\mathbf{X}^{*}]$ for every $\mathbf{X}^{*}\subseteq\operatorname{Form}$ with $\operatorname{Sp}\mathbf{X}^{*}\subsetneq\operatorname{Sp}\mathbf{X}$. Put differently, no strict subspace $V^{\prime}\lneq\operatorname{Sp}\mathbf{X}$ generates $\mathbf{P}$ (where $V^{\prime}$ has a basis of forms $\mathbf{X}^{*}$).
Definition 2.4 (rank-regularity).
For any $t\geq 0$:
  • $\mathbf{X}\in\operatorname{Form}^{r}$ is $t$-rank-regular if $\operatorname{rk}(X)\geq tr$ for all $X\in\operatorname{Sp}\mathbf{X}$ outside of a strict subspace.
  • $\mathbf{P}\in\operatorname{Poly}^{m}(\mathbb{F})$ has a $t$-rank-regular decomposition of size $r$ if $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ is a minimal decomposition for some $t$-rank-regular $\mathbf{X}\in\operatorname{Form}^{r}(\mathbb{F})$.
Note that “all $X\in\operatorname{Sp}\mathbf{X}$ outside of a strict subspace” refers, explicitly, to all $X\in(\operatorname{Sp}\mathbf{X})\setminus U$ for some strict subspace $U\lneq\operatorname{Sp}\mathbf{X}$. [^10] For example, $U$ might be $\operatorname{Sp}(\mathbf{X}^{\prime})$, provided $\mathbf{X}$ is linearly independent. In other words, $\mathbf{X}$ is $t$-rank-regular if $\operatorname{Sp}\{X\in\operatorname{Sp}\mathbf{X}|\operatorname{rk}(X)<tr\}\neq\operatorname{Sp}\mathbf{X}$. This aspect [^11] Equivalently, this is about high-rank pencils: if $V\leq\operatorname{Poly}$ has a pencil $W$ ($V\setminus U$, $U\lneq V$) with $\operatorname{rk}(P)\geq t\dim(V)$ for every $P\in W$, then any tuple of forms spanning $V$ is $t$-rank-regular (and vice versa). of the definition is one reason that we are able to obtain strong bounds for this regularity lemma.
Theorem 2.5 (rank-regularity lemma).
Every $\mathbf{P}\in\operatorname{Poly}^{m}_{d}$ has a $t$-rank-regular decomposition of size at most $((2t+1)dm)^{2^{d}}$.
2.1.1 Linear spaces of polynomials
Here we prove several useful properties of linear spaces of polynomials, which will be used in the proof of the rank-regularity lemma and the weak regularity lemma. We begin with some convenient definitions.
Henceforth, for $V\leq\operatorname{Poly}$ and $R\in\mathbb{N}$, we denote
$U_{R}(V)=\operatorname{Sp}\{X\in V\,|\,\operatorname{rk}(X)<R\}.$
Note that $\mathbf{X}\in\operatorname{Form}^{k}$ is $t$-rank-regular if and only if $U_{R}(V)\neq V$ for $V=\operatorname{Sp}\mathbf{X}$ and $R=tk$.
We say that a linear space $V\leq\operatorname{Poly}$ is homogeneous if $V$ is spanned by forms. Equivalently, $V$ is the sum—and thus the direct sum—of its homogeneous parts: $V=\sum_{i\geq 0}V_{i}$ where $V_{i}:=V\cap\operatorname{Form}_{i}$. We abbreviate $V^{\uparrow}=V_{\deg(V)}$, that is,
$V^{\uparrow}=\{X\in V|\deg(X)=\deg(V)\},$
where we recall the notation $\deg(V)=\max_{X\in V}\deg(X)$.
We will need the following closure property of the subspaces $U_{R}(V)$.
Lemma 2.6.
Let $V\leq\operatorname{Poly}$, $R\in\mathbb{N}$. Every $X\in V$ with $\deg(X)<\deg(U_{R}(V))$ satisfies $X\in U_{R}(V)$.
Proof.
Abbreviate $U=U_{R}(V)$. Observe that there exists $X_{0}\in U^{\uparrow}$ satisfying $\operatorname{rk}(X_{0})\leq R$. We claim that $X_{0}+X\in U$. Indeed, since $\deg(X)<\deg(U)=\deg(X_{0})$, we have $\operatorname{rk}(X_{0}+X)=\operatorname{rk}((X_{0}+X)^{\uparrow})=\operatorname{rk}(X_{0}^{\uparrow})=\operatorname{rk}(X_{0})\leq R$. Thus, as desired, $X\in U$, being the difference of two members of $U$, $X=(X_{0}+X)-X_{0}$. ∎
Lemma 2.7.
Let $V\leq\operatorname{Poly}$, $R\in\mathbb{N}$. If $V$ is homogeneous then so is $U_{R}(V)$.
Proof.
Put $U=U_{R}(V)$. For $X\in\operatorname{Poly}$, denote by $X^{(i)}\in\operatorname{Form}_{i}$ the degree-$i$ portion of $X$ (so that $X=\sum_{i}X_{i}$). First, we prove another closure property of $U$ (in a homogeneous $V$): for every $X\in V$, if $\operatorname{rk}(X)<R$ then $X^{(i)}\in U$ for every $i$. If $i=\deg(X)$ (that is, $X^{(i)}=X^{\uparrow}$) then $\operatorname{rk}(X^{\uparrow})=\operatorname{rk}(X)<R$; since $V$ is homogeneous, $X^{\uparrow}\in V$, so $X^{\uparrow}\in U$, as needed. If $i<\deg(X)$ then $\deg(X^{(i)})<\deg(X)\leq\deg(U)$, since $X\in U$, so by Lemma 2.6, $X^{(i)}\in U$. This proves our claim.
Let $\mathbf{X}=\{X\in V|\operatorname{rk}(X)<R\}$, so that $U=\operatorname{Sp}(\mathbf{X})$ by definition. By the claim above, $\mathbf{X}^{*}:=\{X^{(i)}|X\in\mathbf{X},i\leq\deg(X)\}\subseteq\operatorname{Form}$ satisfies $\mathbf{X}^{*}\subseteq U$. Since $U=\operatorname{Sp}(\mathbf{X})\subseteq\operatorname{Sp}(\mathbf{X}^{*})\subseteq U$, we deduce that $U$ has a spanning set of forms, which completes the proof. ∎
Note that if $U_{R}(V)$ is homogeneous then, in particular,
$U_{R}(V^{\uparrow})=U_{R}(V)\cap V^{\uparrow}.$ (1)
Corollary 2.8.
Let $V\leq\operatorname{Poly}$, $R\in\mathbb{N}$. If $V$ is homogeneous, then $U_{R}(V)=V$ if and only if $U_{R}(V^{\uparrow})=V^{\uparrow}$.
Proof.
First, suppose $U_{R}(V)=V$. Then $\deg(U_{R}(V))=\deg(V)$, and so by Lemma 2.7 and (1), we obtain $U_{R}(V^{\uparrow})=U_{R}(V)\cap V^{\uparrow}=V^{\uparrow}$, as needed.
Next, suppose $U_{R}(V^{\uparrow})=V^{\uparrow}$. Then $\deg(U)=\deg(V)$, since $U_{R}(V)\cap V^{\uparrow}\neq\emptyset$. By Lemma 2.6, $V\setminus V^{\uparrow}\subseteq U_{R}(V)$. Then $V=V^{\uparrow}+(V\setminus V^{\uparrow})\subseteq U_{R}(V)+U_{R}(V)=U_{R}(V)\subseteq V$, that is, $U_{R}(V)=V$, as needed. ∎
2.1.2 Proof of the rank-regularity lemma
Theorem 2.5 will be proved using a repeated application of the following decomposition lemma. It shows that if a tuple of forms $\mathbf{X}$, all of the same degree $d$, is not $t$-rank-regular, then we can decompose it into not-too-many forms of varying degrees, all strictly smaller than $d$. Crucially, the lemma considers not just individual low-rank forms, but the whole subspace they span.
Lemma 2.9.
If $\mathbf{X}\in\operatorname{Form}_{d}^{r}(\mathbb{F})$ is not $t$-rank-regular, then $\mathbf{X}\subseteq\mathbb{F}[\mathbf{Y}]$ for some $\mathbf{Y}\in\operatorname{Form}^{R}$ with $\deg(\mathbf{Y})<d$ and $R\leq 2tr^{2}$.
Proof.
Put $V=\operatorname{Sp}\mathbf{X}$ ($\leq\operatorname{Form}_{d}$) and $k=\dim V$ ($\leq r$). Let
$\mathbf{X^{-}}=\{X\in V\,|\,\operatorname{rk}(X)<tr\}.$
Suppose $\mathbf{X}$ is not $t$-rank-regular, so that $\operatorname{Sp}(\mathbf{X^{-}})=V$. Thus, $\mathbf{X^{-}}$ contains a basis $\mathbf{B}=(B_{1},\ldots,B_{k})\in\operatorname{Form}^{k}_{d}$ for $V$. This has the following implications:
  • Since $\mathbf{X}\subseteq\operatorname{Sp}\mathbf{B}$ (=V=V), there is a linear map $\mathbf{L}\colon\mathbb{F}^{k}\to\mathbb{F}^{r}$ such that $\mathbf{X}=\mathbf{L}(\mathbf{B})$.
  • Since $\mathbf{B}\subseteq\mathbf{X^{-}}\cap\operatorname{Form}$, and by the definition of $\operatorname{rk}$ (recall Definition 2.3 and the subsequent remark), we have $\mathbf{B}=\mathbf{Q}(\mathbf{Y})$ for some quadratic $\mathbf{Q}\in\operatorname{Form}_{2}^{k}$ and some $\mathbf{Y}\in\operatorname{Form}^{R}$ with $\deg(\mathbf{Y})<d$, where $R\leq\sum_{i\in[k]}2\operatorname{rk}(B_{i})<2tr\cdot k$.
Combining the above, we deduce that $\mathbf{X}=\mathbf{L}(\mathbf{Q}(\mathbf{Y}))=\mathbf{F}(\mathbf{Y})$, where $\mathbf{F}:=\mathbf{L}(\mathbf{Q})\in\operatorname{Form}_{2}^{r}$ and $\mathbf{Y}\in\operatorname{Form}^{R}$ with $\deg(\mathbf{Y})<d$ and $R<2tr\cdot k\leq 2tr^{2}$, which completes the proof. ∎
Proof of Theorem 2.5.
Let $\mathbf{X^{0}}\in\operatorname{Form}^{r_{0}}$ be obtained from $\mathbf{P}\in\operatorname{Poly}^{m}$ by replacing each component of $\mathbf{P}$ with its homogeneous parts of degree $1,2,\ldots,d$. Note that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ and $r_{0}\leq dm$.
We claim that for some $s\geq 0$, there are $s$ homogeneous polynomial maps $\mathbf{X^{1}},\ldots,\mathbf{X^{s}}$, with $\mathbf{X^{i}}\in\operatorname{Form}^{r_{i}}$ for $i=1,\ldots,s$, such that $\mathbf{P}\in\mathbb{F}[\mathbf{X^{s}}]$ is a $t$-rank-regular decomposition, and for every $i\geq 0$:
  • $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{i}}]$,
  • $\deg(\mathbf{X^{i}})\leq d-i$, and
  • $r_{i}\leq(2t+1)^{2^{i}-1}(dm)^{2^{i}}$.
[…]
For each $i$ starting with $i=0$, given $\mathbf{X^{i}}$ we proceed as follows:
  • Consider a minimal subspace $W$ of $\operatorname{Sp}(\mathbf{X^{i}})$, with a basis of forms $\mathbf{Y}\in\operatorname{Form}^{r^{\prime}}$, satisfying $\mathbf{P}\subseteq\mathbb{F}[W]$. Then $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]$ is a minimal decomposition, and moreover, $\deg(\mathbf{Y})\leq\deg(W)\leq\deg(\mathbf{X^{i}})$, and $r^{\prime}=\dim W\leq\dim\operatorname{Sp}(\mathbf{X^{i}})\leq r_{i}$.
  • Now, if $\mathbf{P}\in\mathbb{F}[\mathbf{Y}]$ is $t$-rank-regular then set $\mathbf{X^{i}}=\mathbf{Y}$ and $s=i$, finishing the construction (so $\mathbf{X^{s}}\in\operatorname{Form}^{r_{s}}$). Otherwise, $\mathbf{Y}$ is not $t$-rank-regular. We will use $\mathbf{Y}$ to construct $\mathbf{X^{i+1}}$ with $\deg(\mathbf{X^{i+1}})<\deg(\mathbf{X^{i}})$. Recall that in order to apply Lemma 2.9, we need a tuple of forms satisfying two conditions: they are all of the same degree, and the tuple is not $t$-rank-regular.
    Put $\ell=\deg(\mathbf{Y})$, and let $\mathbf{Y^{*}}$ be obtained from $\mathbf{Y}$ by replacing every component polynomial of degree smaller than $\ell$ by an arbitrary component polynomial of degree $\ell$; that is, each component of $\mathbf{Y}$ that lies outside of $\mathbf{Y}^{\uparrow}$ is replaced in $\mathbf{Y^{*}}$ by a polynomial from $\mathbf{Y}^{\uparrow}$. [^13] That is, $\mathbf{Y^{*}}$ is obtained from $\mathbf{Y}^{\uparrow}$ by arbitrarily duplicating components to achieve size $r^{\prime}$. Note that $\mathbf{Y^{*}}\in\operatorname{Form}^{r^{\prime}}_{\ell}$ (whereas $\mathbf{Y}\in\operatorname{Form}^{r^{\prime}}$). We claim that $\mathbf{Y^{*}}$ is not $t$-rank-regular, so that we would be able to apply Lemma 2.9 on it. We have that $\mathbf{Y^{*}}$ and $\mathbf{Y}^{\uparrow}$ are equal as sets: every component of $\mathbf{Y^{*}}$ lies in $\mathbf{Y}^{\uparrow}$, and every component of $\mathbf{Y}^{\uparrow}$ lies in $\mathbf{Y^{*}}$ (though $\mathbf{Y^{*}}$ has the same size as $\mathbf{Y}$). Put $V=\operatorname{Sp}\mathbf{Y}$. Since $\mathbf{Y}$ is not $t$-rank-regular, we have $U_{R}(V)=V$ for $R=tr^{\prime}$. By Corollary 2.8, $U_{R}(V^{\uparrow})=V^{\uparrow}$. This implies that $\mathbf{Y^{*}}$ is not $t$-rank-regular, as claimed, using that $V^{\uparrow}=(\operatorname{Sp}\mathbf{Y})^{\uparrow}=\operatorname{Sp}(\mathbf{Y}^{\uparrow})=\operatorname{Sp}(\mathbf{Y^{*}})$ and $|\mathbf{Y^{*}}|=|\mathbf{Y}|$ (that is, $R=t|\mathbf{Y^{*}}|$).
    Apply Lemma 2.9 on $\mathbf{Y^{*}}$. Merge the resulting tuple of forms with the forms in $\mathbf{Y}$ of degree smaller than $\ell$ to obtain $\mathbf{X^{i+1}}\in\operatorname{Form}^{r_{i+1}}$ with $\mathbf{Y}\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, $\deg(\mathbf{X^{i+1}})<\deg(\mathbf{X^{i}})$, and $r_{i+1}\leq 2tr_{i}^{2}+r_{i}\leq(2t+1)r_{i}^{2}$.
To prove that this construction gives the claim above, we proceed by induction on $i\geq 0$. The induction base $i=0$ follows trivially from the construction of $\mathbf{X^{0}}$, as $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X^{0}}]$ with $\deg(\mathbf{X^{0}})\leq d$, and $r_{0}\leq dm=(2t)^{0}(dm)^{1}$. For the induction step we have, by the construction of $\mathbf{X^{i+1}}$ and the induction hypothesis, that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]\subseteq\mathbb{F}[\mathbf{X^{i+1}}]$, that $\deg(\mathbf{X^{i+1}})\leq\deg(\mathbf{X^{i}})-1\leq d-(i+1)$, and that $r_{i+1}\leq(2t+1)r_{i}^{2}\leq(2t+1)((2t+1)^{2^{i}-1}(dm)^{2^{i}})^{2}=(2t+1)^{2^{i+1}-1}(dm)^{2^{i+1}}$, as needed. This completes the proof. ∎
2.2 Bias and weak regularity
We first define the notion of bias of a polynomial, which measures deviation from equidistribution. Henceforth, we fix any nontrivial additive character $\chi\colon\mathbb{F}\to\mathbb{C}$ (for concreteness, if $\mathbb{F}=\mathbb{F}_{p}$ is a prime field, take $\chi(y)=\omega^{y}$, where $\omega$ is the principle root of unity, $\omega=e^{2\pi i/p}$, and $y$ on the right hand side is viewed as an element of $\{0,\ldots,p-1\}$.) The bias of $P\in\operatorname{Poly}(\mathbb{F})$ is
$\operatorname{bias}(P)=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}\chi(P(x)).$
We note that $|\operatorname{bias}(P)|$ is closely related to the statistical distance between the distribution of $P$ and the uniform distribution on $\mathbb{F}$ (see e.g. [BV10], Claim 33).
Towards proving that rank-regularity implies weak regularity, we will use the best bound known for the structure versus randomness problem for polynomials, which is almost linear (and linear if $|\mathbb{F}|\geq\operatorname{rk}(P)^{\epsilon}$). Here, $\mathcal{L}_{\mathbb{F}}(r):=\log_{|\mathbb{F}|}(r+1)+1$.
Theorem 2.10 ([MZ24], Corollary 8.5141414See also Remark 8.3 for the bound on $c_{d}$.).
For $P\in\operatorname{Poly}_{d}(\mathbb{F})$ with $\operatorname{char}(\mathbb{F})>d$,
$\operatorname{rk}(P)\geq r\quad\Rightarrow\quad|\operatorname{bias}(P)|\leq|\mathbb{F}|^{-c_{d}r/\mathcal{L}_{\mathbb{F}}(r)}$
and $c_{d}=O_{d}(1)$, namely $c_{d}=2^{-d^{1+o(1)}}$.
We also remark that one can replace the use of Theorem 2.10 in our proof by earlier bounds of Milićević [Mil19] or Janzer [Jan20], and still obtain reasonable, albeit somewhat worse bounds in Theorem 2.2.
Using the notion of bias, a Fourier-analytic argument would help us to deduce weak regularity.
Lemma 2.11.
Let $U\lneq V\leq\operatorname{Poly}(\mathbb{F}_{q})$ with $\dim(V)=k$. Suppose $|\operatorname{bias}(X)|\leq\epsilon q^{-k}$ for all $X\in V\setminus U$. Then any basis $\mathbf{X}\in\operatorname{Form}^{k}$ of $V$ that contains a basis of $U$, with $X_{1}\notin U$ and $\deg(X_{1})=\deg(\mathbf{X})$, is weak $\epsilon$-regular.
Recall that $\mathbf{X}$ being weak $\epsilon$-regular implies that for every $y\in\mathbb{F}^{k}$,
$\big|\Pr[\mathbf{X}=\mathbf{y}]-|\mathbb{F}|^{-1}\Pr[\mathbf{X}^{\prime}=\mathbf{y}^{\prime}]\big|\leq\epsilon|\mathbb{F}|^{-k}.$
To prove Lemma 2.11, we will bound the left hand side by $\max_{a_{1}\neq 0}|\operatorname{bias}(a\cdot\mathbf{X})|$.
It will be convenient to use a more geometric formulation of weak $\epsilon$-regularity: for every line $\ell\subseteq\mathbb{F}^{k}$ in the direction of $e_{1}$ (so $\ell=\{te_{1}+\mathbf{z}|t\in\mathbb{F}\}$ for some $\mathbf{z}\in\mathbb{F}^{k}$),
$\forall y\in\ell\colon|\Pr[\mathbf{X}=\mathbf{y}]-|\mathbb{F}|^{-1}\Pr[\mathbf{X}\in\ell]|\leq\epsilon|\mathbb{F}|^{-k}.$ (2)
Proof of Lemma 2.11.
We will use the standard fact that the distribution $(\Pr[\mathbf{X}=y])_{y\in\mathbb{F}^{k}}$ can be expressed in terms of biases, by using simple Fourier analysis. Indeed, since [^16] Indeed, $\sum_{a\in\mathbb{F}^{m}}\chi(a\cdot v)=\sum_{a\in\mathbb{F}^{m}}\prod_{i=1}^{m}\chi(a_{i}v_{i})=\prod_{i=1}^{m}\sum_{a_{i}\in\mathbb{F}}\chi(a_{i}v_{i})=\prod_{i=1}^{m}\delta[v_{i}=0]|\mathbb{F}|=\delta[v=0]|\mathbb{F}|^{m}$.
$\forall v\in\mathbb{F}^{m}\colon\quad\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{m}}\chi(a\cdot v)=\delta[v=0]:=\begin{cases}1&v=0\\ 0&v\neq 0\end{cases}\;,$ (3)
we have that for every $y\in\mathbb{F}^{k}$:
$\Pr_{x}[\mathbf{X}(x)=y]=\operatorname*{\mathbb{E}}_{x}\,\delta[\mathbf{X}(x)-y=0]=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{k}}\chi(a\cdot(\mathbf{X}(x)-y))=\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{k}}\operatorname{bias}(a\cdot(\mathbf{X}-y)).$ (4)
Recall that $\mathbf{X}$ contains a basis for $U$, and $X_{1}\notin U$. We claim that for every $a\in\mathbb{F}^{k}$ with $a_{1}\neq 0$, we have that $a\cdot\mathbf{X}\notin U$. Indeed, write $a\cdot\mathbf{X}=a_{1}X_{1}+X^{\prime}$ with $X^{\prime}\in\operatorname{Sp}(\mathbf{X}^{\prime})$. Our assumption on $\mathbf{X}$ implies that $U\subseteq\operatorname{Sp}(\mathbf{X}^{\prime})$, so if $a\cdot\mathbf{X}\in U$ then $a_{1}X_{1}\in U-X^{\prime}\subseteq\operatorname{Sp}(\mathbf{X}^{\prime})$, a contradiction as $a_{1}\neq 0$. Using our assumption on $U$, we deduce that
$\forall a\in\mathbb{F}^{k}\colon\quad a_{1}\neq 0\quad\Rightarrow\quad|\operatorname{bias}(a\cdot\mathbf{X})|\leq\epsilon|\mathbb{F}|^{-k}.$ (5)
Put $v=e_{1}\in\mathbb{F}^{k}$. Let $\ell\subseteq\mathbb{F}^{k}$ be a line in the direction of $v$, so that $\ell=\{tv+z|t\in\mathbb{F}\}$ for some $z\in\ell$. We claim that the following equality holds:
$\Pr_{x}[\mathbf{X}(x)\in\ell]=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v=0\end{subarray}}\,\operatorname{bias}(a\cdot(\mathbf{X}-z)).$ (6)
Indeed,
$\Pr_{x}[\mathbf{X}(x)\in\ell]$ $=\sum_{t\in\mathbb{F}}\Pr_{x}[\mathbf{X}(x)=tv+z]=\sum_{t\in\mathbb{F}}\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{k}}\operatorname*{\mathbb{E}}_{x}\chi(a\cdot(\mathbf{X}(x)-(tv+z)))$
$=\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{k}}\big(\sum_{t\in\mathbb{F}}\chi(t(a\cdot v))\big)\operatorname*{\mathbb{E}}_{x}\chi(a\cdot(\mathbf{X}(x)-z))$
$=\operatorname*{\mathbb{E}}_{a\in\mathbb{F}^{k}}\big(|\mathbb{F}|\delta[a\cdot v=0]\big)\operatorname{bias}(a\cdot(\mathbf{X}-z))=\frac{1}{|\mathbb{F}|^{k-1}}\sum_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v=0\end{subarray}}\operatorname{bias}(a\cdot(\mathbf{X}-z))$
where (4) is used in the second equality, and (3) is used in the penultimate equality with $m=1$. Combining the point equality (4) and the line equality (6), we deduce (for $y=z$)
$\Pr_{x}[\mathbf{X}(x)=z]=|\mathbb{F}|^{-1}\Pr_{x}[\mathbf{X}(x)\in\ell]+|\mathbb{F}|^{-k}\sum_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v\neq 0\end{subarray}}\,\operatorname{bias}(a\cdot(\mathbf{X}-z)).$
Now, we bound the magnitude of the rightmost term as follows:
$|\mathbb{F}|^{-k}\Big|\sum_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v\neq 0\end{subarray}}\,\operatorname{bias}(a\cdot(\mathbf{X}-z))\Big|\leq\max_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v\neq 0\end{subarray}}|\operatorname{bias}(a\cdot(\mathbf{X}-z))|=\max_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v\neq 0\end{subarray}}|\operatorname{bias}(a\cdot\mathbf{X})|.$
Therefore,
$\Big|\Pr_{x}[\mathbf{X}(x)=y]-|\mathbb{F}|^{-1}\Pr_{x}[\mathbf{X}(x)\in\ell]\Big|\leq\max_{\begin{subarray}{c}a\in\mathbb{F}^{k}\colon\\ a\cdot v\neq 0\end{subarray}}|\operatorname{bias}(a\cdot\mathbf{X})|\leq\epsilon|\mathbb{F}|^{-k},$
using (5) in the last inequality. Thus, $\mathbf{X}$ is weak $\epsilon$-regular, as desired. ∎
2.3 Putting everything together
The proof of the weak regularity lemma will rely on Theorem 2.5, Theorem 2.10, and Lemma 2.11. Careful work will be required to convert rank-regularity, which gives information only about the highest-degree portion (recall Definition 2.3) of the polynomials in our linear space, to weak regularity, which takes into account polynomials of all degrees, while still retaining the homogeneity of the parts. Moreover, rank-regularity depends only on the linear space spanned by the parts, whereas weak regularity depends on the particular parts (and so does not allow the first part to be repeated, for example).
Proof of Theorem 2.2.
Let $t$ be such that the bound in Theorem 2.10 satisfies $q^{-c_{d}tk/\mathcal{L}_{\mathbb{F}_{q}}(tk)}<\epsilon q^{-k}$ for every $k\leq((2t+1)dm)^{2^{d}}$ (recall $\epsilon=q^{-r})$. We observe that $t=2^{d^{1+o(1)}}(r+1)^{1+o(1)}\log(m)$ suffices. [^17] In fact, this choice of $t$ suffices even if the right hand side is $\epsilon^{k}q^{-k}$. If we instead write $\epsilon=q^{-(r-1)}$ for some $r$, then this becomes $q^{-rk}$ for a similar $t$. Apply Theorem 2.5 to find a $t$-rank-regular decomposition of $\mathbf{P}$ of size
$s\leq((2t+1)dm)^{2^{d}}=(2m(r+1))^{2^{d(1+o(1))}},$
so that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]$ is a minimal decomposition and $\mathbf{Y}\in\operatorname{Form}^{s}$ is $t$-rank-regular. To complete the proof, we will use $\mathbf{Y}$ to obtain a weak $\epsilon$-regular partition of $\mathbf{P}$ of size at most $s$.
Put $V=\operatorname{Sp}\mathbf{Y}$ and $k=\dim V$ ($\leq s$). Consider the subspace of $V$,
$U=\operatorname{Sp}\{X\in V\,|\,\operatorname{rk}(X)<tk\},$
and subset of $V$,
$U_{\epsilon}=\{X\in V\,|\,\operatorname{bias}(X)\geq\epsilon q^{-k}\}.$
Then:
  1. 1.
    $U$ is homogeneous. This is by Lemma 2.7 and since $V$ is (it is spanned by $\mathbf{Y}\subseteq\operatorname{Form}$).
  2. 2.
    $V^{\uparrow}\setminus U\neq\emptyset$. Indeed, by (1) we have $U\cap V^{\uparrow}=U_{tk}(V^{\uparrow})$, and by Corollary 2.8 we have $U_{tk}(V^{\uparrow})\neq V^{\uparrow}$, since $U_{tk}(V)\neq V$ (as $\mathbf{Y}$ is $t$-rank-regular).
  3. 3.
    $U_{\epsilon}\subseteq U$. Indeed, for $X\in U_{\epsilon}$ we have $\operatorname{rk}(X)<tk$ by Theorem 2.10 and our choice of $t$.
We conclude that $V$ has a basis of forms $\mathbf{X}\in\operatorname{Form}^{k}$ that contains a basis (of forms) of $U$, by 1, as well as a member of $V^{\uparrow}\setminus U$ by 2 (as $V^{\uparrow}\setminus U\neq\emptyset,\{0\}$). Moreover, $U$ satisfies $(V\setminus U)\cap U_{\epsilon}=\emptyset$ by 3; that is, $|\operatorname{bias}(X)|\leq\epsilon q^{-k}$ for every $X\in V\setminus U$. Therefore, we can apply Lemma 2.11 on $U$, $V$, and $\mathbf{X}$, and deduce that $\mathbf{X}$ is weak $\epsilon$-regular.
Finally, since $\operatorname{Sp}\mathbf{Y}=V=\operatorname{Sp}\mathbf{X}$, we have that $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]=\mathbb{F}[\mathbf{X}]$; moreover, by definition, since $\mathbf{P}\subseteq\mathbb{F}[\mathbf{Y}]$ is a minimal decomposition, so is $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$. In particular, $\mathbf{P}\nsubseteq\mathbb{F}[\mathbf{X}^{\prime}]$, since $\mathbf{X}^{\prime}\subseteq\operatorname{Form}$ satisfies $\operatorname{Sp}\mathbf{X}^{\prime}\subsetneq\operatorname{Sp}\mathbf{X}$. Thus, $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ is a weak $\epsilon$-regular decomposition. The size of this decomposition is $k$ ($\leq s$), which, as discussed above, completes the proof. ∎
3 Applications
In this section we obtained several results along the theme that one can deduce strong bounds on the algebraic structure of polynomials, and polynomial maps, based solely on their image. The proofs use the weak regularity lemma.
We begin with the following standard quantification of algebraic structure, which constraints the degrees of the polynomials in the decomposition.
Definition 3.1 ($t$-rank).
For $t\geq 1$ and $P\in\operatorname{Poly}$, call $t$-reducible if $P$ is the product of polynomials of degree at most $t$. Then
$\operatorname{rk}_{t}(P)=\min\big\{r\,\big|\exists R_{1},\ldots,R_{r}\in\operatorname{Poly}(\mathbb{F})\text{ $t$-reducible}\colon P=R_{1}+\cdots+R_{r}\big\}.$
This definition was used by Karam [Kar23], and is closely related to an earlier definition of Green and Tao [GT09], $\operatorname{rank}_{t}(P)$, which expresses $P$ over a finite field as any function of $\operatorname{rank}_{t}(P)$ polynomials of degree at most $t$. (Note that $\operatorname{rank}_{t}(P)\leq t\operatorname{rk}_{t}(P)$, as each $t$-reducible can be expressed in terms of at most $t$ polynomials; thus, proving an upper bound on $\operatorname{rk}_{t}(P)$, as we will do below, would imply a similar upper bound on $\operatorname{rank}_{t}(P)$.)
Remark 3.2.
If $\operatorname{rk}_{t}(P)$ is small, for $t$ bounded away from $\deg(P)$, then $P$ is “correlated” with a polynomial having a considerably smaller degree (unlike most polynomials of a given degree). More precisely, by a standard Fourier-analytic argument (see e.g. [Jan20]), $P\in\operatorname{Poly}(\mathbb{F}_{q})$ has correlation $|\operatorname{bias}(P-Q)|\geq q^{-\operatorname{rank}_{t}(P)}$ ($\geq q^{-t\operatorname{rk}_{t}(P)}$) with some polynomial $Q$ of degree at most tt. [^18] In the context of the Gowers inverse theorem, this is sometimes denoted $\left\lVert\omega^{P}\right\rVert_{u^{t+1}}\geq q^{-\operatorname{rank}_{t}(P)}$.
One can extend $\operatorname{rk}_{t}$ to polynomial tuples as follows. Define $\operatorname{rk}_{t}(\mathbf{P})$ for $\mathbf{P}\in\operatorname{Poly}^{m}$ to be the least number $r$ of $t$-irreducible polynomials $R_{1},\ldots,R_{r}$ such that $\mathbf{P}\subseteq\operatorname{Sp}\{R_{1},\ldots,R_{r}\}$. Note that for $m=1$, this indeed reduces to Definition 3.1.
An application of the weak regularity lemma that is quite simple to state involves polynomial maps whose image is a line-free set. Recall that a line in $\mathbb{F}_{q}^{m}$ is a set of $q$ elements of the form $\{ta+b\,|\,t\in\mathbb{F}_{q}\}$ for some $a,b\in\mathbb{F}_{q}^{m}$ ($a\neq 0$). Henceforth, we denote
$f(m,d)=(2m)^{2^{d(1+o(1))}}.$
Theorem 3.3.
Let $\mathbf{P}\in\operatorname{Poly}_{d}^{m}(\mathbb{F})$ with $d<\operatorname{char}(\mathbb{F})$. If $\operatorname{Im}\mathbf{P}\subseteq\mathbb{F}^{m}$ does not contain a line then
$\operatorname{rk}_{d/2}(\mathbf{P})\leq f(m,d).$
In particular, the $m=1$ case of Theorem 3.3 is a statement about non-surjective (multivariate) polynomials.
Remark 3.4.
Theorem 3.3 can be viewed as a result about systems of polynomials equations. Note that for any $\mathbf{P}$ consisting of homogeneous polynomials of degree $d$ over $\mathbb{F}_{q}$ with $\gcd(d,q-1)=1$, if the system of equations $\mathbf{P}(x)=y_{0}$ has a solution for some $y_{0}$, then so does $\mathbf{P}(x)=y$ for each $y$ on the line through $y_{0}$ and 0 (indeed, if $\mathbf{P}(x)=y_{0}$ then $\mathbf{P}(cx)=c^{d}\mathbf{P}(x)=ty_{0}$ for $t=c^{d}$). Thus, either $\mathbf{P}(x)=y$ has a solution for each $y$ on some line, or else $\mathbf{P}\equiv 0$. Theorem 3.3 implies that for arbitrary polynomials $\mathbf{P}$ of any degrees, either the system $\mathbf{P}(x)=y$ has a solution for each $y$ on some line, or else $\mathbf{P}$ has a strong algebraic structure (namely, small $\operatorname{rk}_{d/2}(\mathbf{P})$).
Theorem 3.3 is a special case of a theorem about the degree of curves contained in the image of a polynomial map. We next define the univariate degree of a polynomial map $\mathbf{P}$, which, roughly, measures the sparsity of the image of $P$, or how far $P$ is from being surjective—the larger it is, the further $P$ is from being surjective.
Definition 3.5 (Univariate degree).
For $\mathbf{P}\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$,
$\operatorname{udeg}(\mathbf{P})=\min\{\deg(\mathbf{U})\,|\,\mathbf{U}\colon\mathbb{F}\to\mathbb{F}^{m}\text{ non-constant}\colon\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}\}.$
We include a brief list of properties of univariate degree in Section 3.1 below.
Remark 3.6.
The special case of Definition 3.5 where $(n,m,\operatorname{udeg}(P))=(1,1,1)$ is already very interesting. Indeed, note that for any finite field $\mathbb{F}$, a univariate polynomial $P\colon\mathbb{F}\to\mathbb{F}$ satisfies $\operatorname{udeg}(P)=1$ if and only if $P$ is a permutation of $\mathbb{F}$; such polynomials are naturally called permutation polynomials, and have attracted much attention (see e.g. [Hou15] and the references within).
Our main result in this section is as follows.
Theorem 3.7.
For every $\mathbf{P}\in\operatorname{Poly}_{d}^{m}(\mathbb{F})$ with $d<\operatorname{char}(\mathbb{F})$ and with $u=\operatorname{udeg}(\mathbf{P})$, we have
$\operatorname{rk}_{d/u}(\mathbf{P})\leq f(m,d).$
We highlight two extreme special cases. One is where $\operatorname{udeg}(\mathbf{P})$ is very large, namely $\operatorname{udeg}(\mathbf{P})>\deg(\mathbf{P})/2$, in which case we directly obtain a small depth-$3$ (homogeneous) arithmetic formula for $\mathbf{P}$: each component $P_{i}$ is expressed as a linear combination of (the same) few products of degree-$1$ polynomials.
Corollary 3.8.
For $\mathbf{P}\in\operatorname{Poly}_{d}^{m}(\mathbb{F})$ with $d<\operatorname{char}(\mathbb{F})$, if $\operatorname{udeg}(\mathbf{P})>d/2$ then $\operatorname{rk}_{1}(\mathbf{P})\leq f(m,d)$.
The other extreme special case is Theorem 3.3, which is the special case where $\operatorname{udeg}(\mathbf{P})>1$. Indeed, $\operatorname{udeg}(\mathbf{P})=1$ means $\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}$ for some univariate $\mathbf{U}\colon\mathbb{F}\to\mathbb{F}^{m}$ with $\deg(\mathbf{U})=1$, and observe that $\deg(\mathbf{U})=1$ if and only if $\operatorname{Im}\mathbf{U}$ is a line. Indeed, $\deg(\mathbf{U})=1$ means $U_{i}(t)=ta_{i}+b_{i}$ with $a_{i},b_{i}\in\mathbb{F}$, or equivalently, $\mathbf{U}(t)=ta+b$ with $a,b\in\mathbb{F}^{m}$, where $a\neq 0$. Thus, if $\operatorname{Im}\mathbf{P}$ does not contain a line then $\operatorname{udeg}(\mathbf{P})\geq 2$, so Theorem 3.7 implies $\operatorname{rk}_{d/2}(\mathbf{P})\leq f(m,d)$ (that is, Theorem 3.3).
3.1 Properties of univariate degree
Below is a brief list of observation regarding the notion of univariate degree, which we include for the sake of completeness. For simplicity, we state them for the case of a single polynomial ($m=1$).
Proposition 3.9.
For $P\in\operatorname{Poly}(\mathbb{F})$, the univariate degree $\deg(P)$ satisfies the following properties:
  1. 1.
    $\operatorname{udeg}(P)\leq\min_{i}\deg_{i}(P)$ ($\leq\deg(P)$).
  2. 2.
    $\operatorname{udeg}(P)<|\mathbb{F}|$.
  3. 3.
    $\operatorname{udeg}(F(\mathbf{X}))\geq\operatorname{udeg}(F)$, for any $F\in\operatorname{Poly}$, with equality if $\mathbf{X}$ is surjective.
  4. 4.
    $\operatorname{udeg}(U)\leq\deg(U)$ for any univariate $U$.
  5. 5.
    $\operatorname{udeg}(aP+b)=\operatorname{udeg}(P)$ for any $a,b\in\mathbb{F}$ with $a\neq 0$.
  6. 6.
    $\operatorname{udeg}(P^{2})\geq 2$ if $\operatorname{char}(\mathbb{F})>2$, with equality if $P$ is surjective.
Proof.
In order:
  1. 1.
    Fix all but the $i$-ith variable to constants so that the resulting univariate polynomial is nonzero. (It is indeed clear that such an assignment exists: we may assume that all variables in $P$ have degree smaller than $|\mathbb{F}|$. Then, $P=\sum_{j}x_{i}^{j}C_{j}$, with some $C_{j}$ a nonzero polynomial and thus also a nonzero function on $\mathbb{F}^{n-1}$.)
  2. 2.
    In the definition of univariate degree, the condition on the univariate $U$ depends only on the function it computes, and $x^{|\mathbb{F}|}=x$ as functions.
  3. 3.
    For $U$ witnessing $\operatorname{udeg}(F(\mathbf{X}))$, we have $\operatorname{Im}(U)\subseteq\operatorname{Im}(F(\mathbf{X}))\subseteq\operatorname{Im}(F)$, so $\operatorname{udeg}(F)\leq\deg(U)=\operatorname{udeg}(F(\mathbf{X}))$. Moreover, if $\mathbf{X}$ is surjective then $\operatorname{Im}(F)=\operatorname{Im}(F(\mathbf{X}))$, and $\operatorname{udeg}(P)$ depends only on $\operatorname{Im}(P)$.
  4. 4.
    By definition. Note that the inequality can be strict (e.g., a permutation polynomial of degree at least $2$).
  5. 5.
    If a univariate $U$ satisfies $\operatorname{Im}(U)\subseteq\operatorname{Im}(P)$, then $\operatorname{Im}(aU+b)\subseteq\operatorname{Im}(aP+b)$, so $\operatorname{udeg}(aP+b)\leq\deg(aU+b)=\deg(U)$, which shows $\operatorname{udeg}(aP+b)\leq\operatorname{udeg}(P)$ if we choose $U$ that witnesses $\operatorname{udeg}(P)$. The equality follows by symmetry.
  6. 6.
    $\operatorname{udeg}(P^{2})\geq\operatorname{udeg}(x^{2})$, and we claim that $\operatorname{udeg}(x^{2})=2$. It is not surjective as $\gcd(2,|\mathbb{F}|-1)=2$, using that $|\mathbb{F}|$ is odd. Thus, $2\leq\operatorname{udeg}(x^{2})\leq\deg(x^{2})=2$. Moreover, if $P$ is surjective then $\operatorname{udeg}(P^{2})=\operatorname{udeg}(x^{2})$.
3.2 Proofs of the applications
First, we will need the following immediate property of weak $\epsilon$-regular tuples. Note that for an arbitrary map $\mathbf{X}\colon\mathbb{F}^{n}\to\mathbb{F}^{r}$, if a line $\ell\subseteq\mathbb{F}^{r}$ in a given direction is chosen uniformly at random, then $\operatorname*{\mathbb{E}}_{\ell}\Pr[\mathbf{X}\in\ell]=q^{-(r-1)}$. We show that if $\mathbf{X}$ is weak regular, then lines that are somewhat popular (that is, close to being as popular as a random line) must be fully contained in the image of $\mathbf{X}$.
Lemma 3.10.
Let $\mathbf{X}\subseteq\operatorname{Form}^{r}(\mathbb{F}_{q})$ be weak $\epsilon$-regular. For every line $\ell\subseteq\mathbb{F}^{r}$ in direction $e_{1}$, if $\Pr[\mathbf{X}\in\ell]>\epsilon q^{-(r-1)}$ then $\ell\subseteq\operatorname{Im}\mathbf{X}$.
Proof.
By the definition of weak $\epsilon$-regularity (recall (2)), for every $y\in\ell$ we have
$\Pr[\mathbf{X}=y]\geq\Pr[\mathbf{X}\in\ell]/q-\epsilon q^{-r}>0.$
Thus $y\in\operatorname{Im}\mathbf{X}$, which completes the proof. ∎
We say that a polynomial decomposition $P=F(\mathbf{X})$ is irredundant if there is no polynomial $F^{\prime}$ with $\operatorname{Supp}F^{\prime}\subsetneq\operatorname{Supp}F$ such that $F^{\prime}(\mathbf{X})=F(\mathbf{X})$. [^20] Recall that $\operatorname{Supp}F$ denotes the support of $F$, meaning the set of monomials appearing with a nonzero coefficient in $F$.
We will need a few more auxiliary results. First, we shows that a decomposition of a polynomial $P$ into forms need not involve terms of higher degree than that of $P$.
Lemma 3.11 (irredundant degree).
For $P\in\operatorname{Poly}$, let $P=F(\mathbf{X})$, with $\mathbf{X}\subseteq\operatorname{Form}$, be an irredundant decomposition. For every monomial $M$ in the support of $F$, $\deg M(\mathbf{X})\leq\deg P$.
Proof.
Since every $X_{i}$ is homogeneous, for every monomial $M\in\operatorname{Supp}(F)$, all the monomials that result from expanding $M(\mathbf{X})$ are of the same degree. Thus, for each $k$, the degree-$k$ part of $F(\mathbf{X})$ is equal to $\sum M(\mathbf{X})$ with the sum over all monomials $M\in\operatorname{Supp}(F)$ with $\deg M(\mathbf{X})=k$. Therefore, if $F^{\prime}$ is obtained from $F$ by removing from its support all monomials $M$ with $\deg M(\mathbf{X})=k$, for $k>\deg F(\mathbf{X})$, then we still have $F^{\prime}(\mathbf{X})=F(\mathbf{X})$. We deduce that if there exists a monomial $M\in\operatorname{Supp}(F)$ with $\deg M(\mathbf{X})>\deg F(\mathbf{X})$, then $P=F(\mathbf{X})$ is not an irredundant decomposition, completing the proof. ∎
We next prove, using the Schwartz-Zippel lemma, that in the image of any polynomial map $\mathbf{X}$ one can find a popular line that avoids a given low-degree zero set.
Lemma 3.12 (Schwartz-Zippel for decompositions).
Let $P\in\operatorname{Poly}_{d}(\mathbb{F}_{q})$ be nonzero. If $P=F(\mathbf{X})$ with $\mathbf{X}\in\operatorname{Poly}^{k}$ (and $F\in\operatorname{Poly}$), then there exists $y\in\mathbb{F}_{q}^{k}$ such that $F(y)\neq 0$ and
$(\Pr[\mathbf{X}=y]\geq)\,\Pr[F\neq 0]\cdot\Pr[\mathbf{X}=y]\geq(1-d/q)q^{-k}.$
Proof.
Consider $\{y\in\mathbb{F}^{k}\,|\,F(y)\neq 0\}$, the complement of the zero set $\operatorname{Z}(F)$ of $F$. We have
$\Pr_{x}[P(x)\neq 0]=\Pr_{x}[\mathbf{X}\notin\operatorname{Z}(F)]=\sum_{y\notin\operatorname{Z}(F)}\Pr[\mathbf{X}=y]\leq q^{k}\Pr[F\neq 0]\max_{y\notin\operatorname{Z}(F)}\Pr[\mathbf{X}=y].$
On the other hand, as $P$ is nonzero, the Schwartz-Zippel lemma gives $\Pr[P\neq 0]\geq 1-d/q$. Combining the above, we get $\Pr[F\neq 0]\max_{y\notin\operatorname{Z}(F)}\Pr[\mathbf{X}=y]\geq(1-d/q)q^{-k}$, as desired. ∎
We combine the above lemmas to find a popular line that the decomposition nontrivially depends on. Denote by $\deg_{1}(F)$ the degree of $F\in\mathbb{F}[x_{1},\ldots,x_{r}]$ in its first variable (that is, the largest exponent of $x_{1}$ in the monomials in the support of $F$).
Corollary 3.13.
Let $P\in\operatorname{Poly}_{d}(\mathbb{F}_{q})$. If $P=F(\mathbf{X})$ with $\mathbf{X}\in\operatorname{Form}^{r}$ is an irredundant decomposition with $\deg_{1}F>0$, then there exists a line $\ell\subseteq\mathbb{F}_{q}^{r}$ in direction $e_{1}$ such that $|F(\ell)|\neq 1$ and $\Pr[\mathbf{X}\in\ell]>(1-d/q)q^{-(r-1)}$.
Proof.
Put $\mathbb{F}=\mathbb{F}_{q}$. Write $F(t,\mathbf{y})=\sum_{i\geq 0}t^{i}C_{i}(y)$ with $C_{i}\in\operatorname{Poly}$, the unique representation of $F$ as a linear combination of the monomials in its support. We claim that the following holds for some $j\geq 1$:
  • $C_{j}(\mathbf{X}^{\prime})$ is nonzero; indeed, otherwise $P=F(\mathbf{X})=F(X_{1},\mathbf{X}^{\prime})=C_{0}(\mathbf{X}^{\prime})$; moreover, $\operatorname{Supp}C_{0}\subseteq\operatorname{Supp}F$, and $\operatorname{Supp}C_{0}\neq\operatorname{Supp}F$ by the assumption $\deg_{1}(F)>0$. This contradicts the assumption that $P=F(\mathbf{X})$ is an irredundant decomposition.
  • $C_{j}(\mathbf{X}^{\prime})$ is of degree smaller than $d$; indeed, since $\mathbf{X}\subseteq\operatorname{Form}$ and the decomposition $P=F(\mathbf{X})$ is irredundant, Lemma 3.11 implies that $\deg(C_{j}(\mathbf{X}^{\prime}))<\deg(X_{1}^{j}C_{j}(\mathbf{X}^{\prime}))\leq\deg(P)\leq d$.
Apply Lemma 3.12 on the nonzero $C_{j}(\mathbf{X}^{\prime})\in\operatorname{Poly}_{d-1}$ to obtain $y_{0}\in\mathbb{F}_{q}^{r-1}$ such that $C_{j}(y_{0})\neq 0$ and $\Pr[\mathbf{X}^{\prime}=y_{0}]>(1-d/q)q^{-(r-1)}$. Observe that $\deg(F(t,y_{0}))\geq j$ ($\geq 1$), so in particular, $F(t,y_{0})$ is a non-constant univariate polynomial. Consider the line $\ell=\{(t,y_{0})\,|\,t\in\mathbb{F}\}\subseteq\mathbb{F}^{r}$. Then $F(\ell)=\{F(t,y_{0})\,|\,t\in\mathbb{F}\}$ is not a singleton. Moreover, observe that $\Pr[\mathbf{X}\in\ell]=\Pr[\mathbf{X}^{\prime}=y_{0}]$. This completes the proof. ∎
We can now prove our main result here.
Proof of Theorem 3.7.
Put $\mathbb{F}=\mathbb{F}_{q}$. Assume $\mathbf{P}$ is nonzero, as otherwise $\operatorname{rk}_{t}(0)=0$ and we are done. Put $\epsilon=1-d/q$. By Theorem 2.2, $\mathbf{P}$ has a weak $\epsilon$-regular decomposition $\mathbf{P}=\mathbf{F}(\mathbf{X})$, with $\mathbf{X}\in\operatorname{Form}^{r}$, of size $r\leq(2m)^{2^{d(1+o(1))}}$. We will prove that $\deg(\mathbf{X})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P})$.
We assume without loss of generality that each decomposition $P_{i}=F_{i}(\mathbf{X})$ is irredundant (by removing monomials from the support of $F_{j}$ if necessary). Since $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, there is a component $P_{j}$ of $\mathbf{P}$ such that $P_{j}\in\mathbb{F}[\mathbf{X}]\setminus\mathbb{F}[\mathbf{X}^{\prime}]$, and therefore $\deg_{1}(F_{j})>0$. Apply Corollary 3.13 on $P_{j}\in\operatorname{Poly}_{d}(\mathbb{F})$ to obtain a line $\ell\subseteq\mathbb{F}^{r}$ in direction $e_{1}$ such that $|F_{j}(\ell)|\neq 1$ and $\Pr[\mathbf{X}\in\ell]>(1-d/q)q^{-(r-1)}$. Now apply Lemma 3.10 on $\ell$, using the weak $\epsilon$-regularity of $\mathbf{X}$ and our choice of $\epsilon$, to deduce that
$\ell\subseteq\operatorname{Im}\mathbf{X}\text{ and }|\mathbf{F}(\ell)|\neq 1.$ (7)
Now, let $\mathbf{U}\colon\mathbb{F}\to\mathbb{F}$ be the curve $\mathbf{U}=\mathbf{F}(\mathbf{L})$, where $\mathbf{L}\colon\mathbb{F}\to\mathbb{F}^{r}$ satisfies $\operatorname{Im}(\mathbf{L})=\ell$. (Explicitly, if $\ell=\{te_{1}+y|t\in\mathbb{F}\}$, then $\mathbf{L}(t)=te_{1}+y$, so $\mathbf{U}(t)=\mathbf{F}(\mathbf{L}(t))=F(t+y_{1},y_{2},\ldots,y_{r})$.) Then $\operatorname{Im}\mathbf{U}=\mathbf{F}(\operatorname{Im}(\mathbf{L}))=\mathbf{F}(\ell)$, which implies that
$\operatorname{Im}\mathbf{U}\subseteq\operatorname{Im}\mathbf{P}\text{ and }\mathbf{U}\text{ is nonconstant},$
since $\mathbf{F}(\ell)\subseteq\mathbf{F}(\operatorname{Im}\mathbf{X})=\operatorname{Im}\mathbf{P}$ and $|\mathbf{F}(\ell)|\neq 1$, by (7). By definition, this means $\operatorname{udeg}(\mathbf{P})\leq\deg(\mathbf{U})$.
Let $i\in[m]$; abbreviate $P=P_{i}$, $F=F_{i}$, and $U=U_{i}$, so that $P=F(\mathbf{X})$ and $U=F(\mathbf{L})$. We claim that the following chain of inequalities holds:
$\deg(U)\leq\deg_{1}(F)\leq\deg(P)/\deg(X_{1})=\deg(P)/\deg(\mathbf{X}).$
The first inequality is immediate from the construction of $U$. The second inequality follows using Lemma 3.11: as $P=F(\mathbf{X})$, for every monomial $M=\prod_{j=1}^{r}x_{j}^{m_{j}(M)}$ in the support of $F$, we have $m_{1}(M)\deg(X_{1})\leq\deg(M(\mathbf{X}))\leq\deg(P)$; since $\deg_{1}(F)=\max_{M}m_{1}(M)$ over $M\in\operatorname{Supp}F$, we indeed obtain that $\deg_{1}(F)\deg(X_{1})\leq\deg(P)$. The final equality follows from the construction of $\mathbf{X}$. We obtain the following degree bound for $\mathbf{X}$, using $i$ satisfying $\deg(U)=\deg(\mathbf{U})$:
$\deg(\mathbf{X})\leq\deg(P)/\deg(U)\leq\deg(\mathbf{P})/\deg(\mathbf{U})\leq\deg(\mathbf{P})/\operatorname{udeg}(\mathbf{P}).$
Finally, as every component $P_{i}$ of $\mathbf{P}$ satisfies $P_{i}=F_{i}(\mathbf{X})$ with $F_{i}\colon\mathbb{F}^{r}\to\mathbb{F}$ and $\deg(F_{i})\leq\deg(P_{i})\leq d$, by Lemma 3.11, we deduce that $P_{1},\ldots,P_{m}$ lie in the span of all products of at most $d$ members of $\mathbf{X}$. Thus, as needed,
$\operatorname{rk}_{d/u}(\mathbf{P})\leq\binom{r+d}{d}=\sum_{i=0}^{d}\binom{r+i-1}{i}\leq\sum_{i=0}^{d}r^{i}\leq(2m)^{2^{d(1+o(1))}}.\qed$
Theorem 1.2, like Theorem 3.7, follows from the weak regularity lemma.
Proof of Theorem 1.2.
Let $P\subseteq\mathbb{F}[\mathbf{P}]$ with $\operatorname{udeg}(\mathbf{P})=u$. As in the proof of Theorem 3.7, using the weak regularity lemma we obtain a decomposition $\mathbf{P}\subseteq\mathbb{F}[\mathbf{X}]$ with $\mathbf{X}\in\operatorname{Form}^{s}$ of size $s\leq(2m)^{2^{d(1+o(1))}}$ such that $\deg(\mathbf{X})\leq d/u$. Observe that $P\in\mathbb{F}[\mathbf{P}]\subseteq\mathbb{F}[\mathbf{X}]$. Using Lemma 3.11, it follows that we can write
$P=\sum_{i=1}^{r}c_{i}\prod_{j}Q_{i,j}.$ (8)
with $Q_{i,j}\in\mathbf{X}$, $c_{i}\in\mathbb{F}$, $\deg(\prod_{j}Q_{i,j})\leq d$, and $r\leq\sum_{i=0}^{d}s^{i}\leq(2m)^{2^{d(1+o(1))}}$. Note that $\deg(Q_{i,j})\leq\deg(\mathbf{X})\leq d/u$.
We claim that we can guarantee that the number of multiplicands in each product $\prod_{j}Q_{i,j}$ is at most $2u$. Note that we can express each product as a product of homogeneous polynomials $\prod_{j=1}^{a_{i}}Q^{\prime}_{i,j}$ all of whose degrees, except at most one of them, satisfy $\frac{1}{2}(d/u)<\deg(Q^{\prime}_{i,j})\leq d/u$; indeed, this follows by repeatedly using the fact that if $\deg(Q_{i,j_{1}}),\deg(Q_{i,j_{2}})\leq\frac{1}{2}(d/u)$ then $\deg(Q_{i,j_{1}}Q_{i,j_{2}})\leq d/u$. Thus, $d\geq\sum_{j=1}^{a_{i}}\deg(Q^{\prime}_{i,j})>(a_{i}-1)\frac{1}{2}(d/u)$, so $a_{i}\leq 2u$ as claimed. Thus, (8) gives a $\sum^{[r]}\prod^{[2u]}\sum\prod^{[d/u]}$-formula for $P$, as needed.
Finally, if $P$ is homogeneous then, since every member of $\mathbf{X}$ is homogeneous, without loss of generality the products in the formula satisfy $\deg(\prod_{j}Q_{i,j})=\deg(P)$ for every $i$, and therefore the formula is homogeneous, as desired. ∎